OpenStudy (anonymous):

please help. Simplify the rational expression. State any excluded values. x^3-36/42 x-7x^2

OpenStudy (anonymous):

\[\frac{x^3-36}{42x-7x^2}\] ?

OpenStudy (anonymous):

thats the problem there wasn't meant to be a space between 42 and x

OpenStudy (anonymous):

i am going to make a bet that it is \[\frac{x^2-36}{42x-7x^2}=\frac{(x-6)(x+6)}{7x(6-x)}\]\] if the numerator is in fact \(x^3-36\) there is not much you can do

OpenStudy (anonymous):

if it is the second one i wrote you can factor and cancel since \(\frac{x-6}{6-x}=-1\)